# FinWorld: An All-in-One Open-Source Platform for End-to-End Financial AI Research and Deployment

### v1

@@ 1. Introduction
Financial AI has revolutionized how we approach market analysis, trading strategies, and investment decisions. From traditional quantitative machine learning (ML) models (Fama and French, 2015; Ke et al., 2017) to modern deep learning (DL) and reinforcement learning (RL) architectures (Jiang et al., 2017; Yu et al., 2019; Liu et al., 2020a, b), the field has witnessed remarkable progress across diverse domains including time series forecasting (Feng et al., 2023; Ding et al., 2020), algorithmic trading (Liu et al., 2020b; Qin et al., 2024), portfolio management (Zhang et al., 2024a; Kim, 2025), and natural language analysis for financial documents (Liu et al., 2023, 2025). However, this rapid advancement has created a fragmented ecosystem where researchers and practitioners must navigate multiple specialized tools, each optimized for specific tasks but lacking seamless integration.
Current financial AI platforms, while valuable in their respective domains, face several critical limitations. For instance, TradeMaster (Sun et al., 2023) provides RL methods for four financial trading tasks but lacks compatibility with traditional quantitative models and financial LLMs. Qlib (Yang et al., 2020) provides robust quantitative investment tools but offers limited support for modern AI paradigms such as LLMs and RL. Similarly, FinRL-Meta (Liu et al., 2022) focuses specifically on RL for trading but suffers from a rigid framework architecture that makes it difficult to integrate new algorithms and lacks support for DL and LLM-based approaches. Most existing platforms are designed for narrow task scopes: some excel at time series forecasting but lack trading simulation, while others offer robust algorithmic trading or portfolio management yet have limited support for modern LLMs and agent paradigms. Furthermore, the integration of heterogeneous financial data ranging from structured market data to unstructured news and reports remains challenging, often requiring complex preprocessing pipelines and custom development efforts.
@@ 2.1. ML, DL, and RL in Financial AI
Financial time series forecasting has evolved from traditional statistical methods like ARIMA (Box et al., 1976) and GARCH (Engle, 1982) to modern DL approaches. ML methods such as LightGBM (Ke et al., 2017) and XGBoost (Chen and Guestrin, 2016) demonstrated superior performance in capturing complex financial patterns, while DL models including LSTM (Hochreiter and Schmidhuber, 1997) and Transformer (Vaswani et al., 2017) architectures have revolutionized temporal modeling. Recent advances in financial forecasting include TimesNet (Wu et al., 2022), DLinear (Zeng et al., 2023a), and Timexer (Liu et al., 2024), which leverage attention mechanisms for improved prediction accuracy. In algorithmic trading, RL methods such as PPO (Schulman et al., 2017) and DQN (Mnih et al., 2015) have enabled adaptive trading strategies (Jiang et al., 2017; Liu et al., 2020b), with recent work exploring transformer-based approaches (Qin et al., 2024). Portfolio management has similarly progressed from classical mean-variance optimization (Markowitz, 1952) to data-driven approaches using ML (Heaton et al., 2017) and RL (Zhang et al., 2024a; Kim, 2025). Despite these advances, existing approaches often focus on individual tasks, lacking unified frameworks for comprehensive financial AI development.
@@ 2.2. LLMs and LLM Agents in Financial AI
The integration of LLMs into financial decision-making has followed two main paths. The traditional approach uses pre-training and supervised fine-tuning (SFT), with models like FinBERT (Araci, 2019) and FLANG (Shah et al., 2022) excelling in financial text understanding. Recent progress includes FinGPT (Liu et al., 2023), BloombergGPT (Wu et al., 2023) with domain-adapted tokenization, and FinQA (Chen et al., 2021) featuring numerical reasoning. A newer direction combines pre-training with reinforcement learning (RL) to enhance reasoning, as shown by Fin-R1 (Liu et al., 2025) and Fino1 (Qian et al., 2025). Despite their strengths, LLM-based methods often lack sequential decision-making, are computationally expensive, and struggle with non-stationary markets.
To overcome the limitations of traditional LLMs in sequential decision-making, researchers have introduced agentic mechanisms such as memory, tool use, and self-reflection. Early systems like FinMem (Yu et al., 2024a) and FinRobot (Zhou et al., 2024) rely on text-only LLMs enhanced with layered memory, profiling, and chain-of-thought reasoning to improve single-asset trading performance. Building on this, FinAgent (Zhang et al., 2024b) represents the first multimodal trading agent, which jointly processes news, structured price data, and K-line chart visuals via a tool-augmented dual-reflection architecture, achieving significant improvements in profit. Recent efforts shift toward collaborative multi-agent systems: FinCon (Yu et al., 2024b) adopts a manager–analyst hierarchy with verbal reinforcement, while TradingAgents (Xiao et al., 2024) models a full trading firm with specialized agents coordinating to enhance returns and risk control. Beyond these architectural innovations, recent work has also explored using RL to finetune LLMs Agents (Feng et al., 2025; Ouyang et al., 2025), enabling dynamic exploration and reasoning capabilities in real-world environments.
@@ 2.3. Financial AI Platforms
The development of financial AI platforms has been driven by the need to provide standardized environments for research and deployment. Qlib (Yang et al., 2020) provides a quantitative investment platform that integrates traditional ML pipelines with financial data processing, offering tools for feature engineering, model training, and portfolio optimization.TradeMaster (Sun et al., 2023) offers a comprehensive RL-based trading framework with modular components for strategy development, backtesting, and evaluation across multiple financial markets. FinRL-Meta (Liu et al., 2022) focuses on reinforcement learning environments and benchmarks, providing standardized market simulators and evaluation metrics for RL-based trading strategies. These platforms have significantly contributed to the advancement of financial AI research by providing researchers and practitioners with accessible tools and standardized evaluation frameworks. However, most existing platforms primarily focus on traditional ML, DL, and RL methods, with limited support for LLMs and LLMs Agents, and lack comprehensive multimodal data integration capabilities. This motivates our proposal of FinWorld, a comprehensive framework that covers most critical financial models and tasks.
@@ 3.5. Reinforcement Learning for LLMs
Group Relative Policy Optimization (GRPO) (Shao et al., 2024) is a policy optimization algorithm designed to efficiently train LLMs in both language and environment-interactive settings. In the context of financial AI, GRPO can be leveraged to endow LLMs with either general financial knowledge and reasoning ability, and specialized trading skills. For example, when applied to financial document analysis or financial question answering, GRPO enables the LLM to align with environment feedback and develop robust understanding of domain-specific content. In these tasks, the LLM is trained to produce high-quality financial responses, with group-level reward normalization providing stable learning signals for fine-tuning.
@@ 4.2. Configuration Layer
The configuration layer of FinWorld is built on mmengine (Contributors, 2022) and provides a unified and extensible system using dictionaries. It centralizes all experimental settings, including datasets, models, training, and evaluation, in a readable and modular format. Support for configuration inheritance and overrides ensures reproducibility, flexibility, and collaborative development. In the meantime, FinWorld uses a registry mechanism to manage core components such as the dataset and the environment. Each class is registered by its type and instantiated from the configuration, allowing for flexible component management and a decoupled system architecture.
@@ 4.3. Dataset Layer
Downloader Module. This module consists of two main components: i) Market Data Downloader, which standardizes access to heterogeneous financial market data sources, such as FMP (Team, 2025b) and Alpaca (Team, 2025a), supporting multiple time resolutions (e.g., daily, minute-level) and data types (e.g., OHLCV, news); and ii) LLM Reasoning Dataset Downloader, which provides access to financial reasoning datasets including FinQA, professional certification materials (e.g., ACCA, CFA), and other financial knowledge bases. The unified downloader abstraction streamlines data acquisition from various APIs and databases, ensuring consistency and scalability across research tasks.
Processor Module. Building on the standardized data access, this module enables users to configure essential preprocessing steps, such as factor computation (e.g., alpha158 (Yang et al., 2020)), feature selection, and normalization, LLM reasoning data processing, and other preprocessing steps according to specific research needs. The processor abstraction facilitates reproducible and customizable data transformations for downstream modeling.
@@ 4.4. Model Layer
DL Models. This module organizes neural architectures into components, layers, and model networks. Components include data embedding for financial time series (Wu et al., 2022), patch embedding (Dosovitskiy et al., 2020), position encodings, attention, transformer blocks, and activation functions. Layers compose these into an embedding layer, an encoder, and a decoder with consistent interfaces. Model networks such as Autoformer, VAE, and GPT-style decoders are instantiated by wiring the layers under a common specification, independent of training and inference.
@@
LLMs Models. This module centralizes access to both proprietary and open-source LLMs via a unified interface. It supports seamless switching among commercial providers (e.g., GPT-4.1, Claude-4-Sonnet, Gemini-2.5-Pro) and efficient local models (e.g., Qwen2.5, Qwen3), with standardized controls for context length, sampling, and cost/latency tracking. Built-in tools include function calling, tool use, and retrieval-augmented generation for financial documents. The module integrates naturally with agents and downstream models, enabling document understanding, information extraction, instruction following.
@@ 5.2. Time Series Forecasting
Methods. We evaluate several ML-based and DL-based time series forecasting models, including: i) ML-based methods: LightGBM (Yang et al., 2020), XGBoost (Yang et al., 2020); ii) DL-based methods: Autoformer (Wu et al., 2021), Crossformer (Zhang and Yan, 2023), ETSformer (Woo et al., 2022), DLinear (Zeng et al., 2023b), TimesNet (Wu et al., 2022), PatchTST (Nie et al., 2022), TimeMixer (Wang et al., 2024b), and TimeXer (Wang et al., 2024a). These models are selected based on their effectiveness in capturing complex patterns and their ability to handle high-dimensional data.
@@ 5.3. Algorithmic Trading
Methods. We systematically evaluate several Rule-based, ML-based, DL-based, and RL-based methods, including i) Rule-based methods: BUY&HOLD, MACD; ii) ML-based Methods: LightGBM, XGBoost; iii) DL-based Methods: Transformer (Vaswani et al., 2017), LSTM, DLinear (Zeng et al., 2023b); iv) RL-based Methods: PPO (Schulman et al., 2017), SAC (Haarnoja et al., 2018).
@@ 5.4. Portfolio Management
[…]
@@ 5.5. LLMs Applications
Methods. For financial reasoning, we evaluate our proposed models FinReasoner against open-source LLMs on the test dataset, including DeepSeek-R1 (Guo et al., 2025), Qwen3-8B (Yang et al., 2025), Fin-R1-7B (Guo et al., 2025), and Qwen2.5-7B-Instruct (Team, 2025d). For trading abilities, we evaluate our own proposed FinReasoner with DeepSeek-R1 (Guo et al., 2025), Qwen3-8B (Yang et al., 2025), Fin-R1-7B (Guo et al., 2025), Qwen2.5-7B-Instruct (Team, 2025d), GPT-4.1, Claude-4-Sonnet, all serving as the backbone LLMs within FinAgent framework.
@@ A.1. Market Dataset
The data processing pipeline includes comprehensive preprocessing steps to ensure data quality and usability. For price data, we implement filtering procedures to remove non-trading days and invalid data points, followed by the calculation of technical indicators using the Alpha158 (Yang et al., 2020) (Table 5) framework, which provides a solid foundation for deep learning modeling. For news text data, we employ a two-stage processing approach: first, we use web scraping scripts to extract raw content from original news links, then we utilize Qwen3-32B to generate concise summaries, enabling efficient text analysis and feature extraction for financial AI applications.
@@
kmid $(close-open)/close$ K-Mid price ratio
kmid2 $(close-open)/(high-low)$ Normalized K-Mid price ratio
klen $(high-low)/open$ K-Length price ratio
kup $(high-max(open,close))/open$ K-Up price ratio
kup2 $(high-max(open,close))/(high-low)$ Normalized K-Up price ratio
klow $(min(open,close)-low)/open$ K-Low price ratio
klow2 $(min(open,close)-low)/(high-low)$ Normalized K-Low price ratio
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
$w$ $|close/close.shift(1)-1|.rolling(w).std()/|close/close.shift(1)-1|.rolling(w).mean()$ Weighted volume moving average
[…]

### v2

@@ 1. Introduction
Financial AI has revolutionized how we approach market analysis, trading strategies, and investment decisions. From traditional quantitative machine learning (ML) models (fama2015five; ke2017lightgbm) to modern deep learning (DL) and reinforcement learning (RL) architectures (jiang2017deep; yu2019model; liu2020adaptive; liu2020finrl), the field has witnessed remarkable progress across diverse domains including time series forecasting (feng2023multi; ding2020hierarchical), algorithmic trading (liu2020finrl; qin2024earnhft), portfolio management (zhang2024reinforcement; kim2025semi), and natural language analysis for financial documents (liu2023fingpt; liu2025fin). However, this rapid advancement has created a fragmented ecosystem where researchers and practitioners must navigate multiple specialized tools, each optimized for specific tasks but lacking seamless integration.
Current financial AI platforms, while valuable in their respective domains, face several critical limitations. For instance, TradeMaster (sun2023trademaster) provides RL methods for four financial trading tasks but lacks compatibility with traditional quantitative models and financial LLMs. Qlib (yang2020qlib) provides robust quantitative investment tools but offers limited support for modern AI paradigms such as LLMs and RL. Similarly, FinRL-Meta (liu2022finrl) focuses specifically on RL for trading but suffers from a rigid framework architecture that makes it difficult to integrate new algorithms and lacks support for DL and LLM-based approaches. Most existing platforms are designed for narrow task scopes: some excel at time series forecasting but lack trading simulation, while others offer robust algorithmic trading or portfolio management yet have limited support for modern LLMs and agent paradigms. Furthermore, the integration of heterogeneous financial data ranging from structured market data to unstructured news and reports remains challenging, often requiring complex preprocessing pipelines and custom development efforts.
@@ 2.1. ML, DL, and RL in Financial AI
Financial time series forecasting has evolved from traditional statistical methods like ARIMA (box1976time) and GARCH (engle1982autoregressive) to modern DL approaches. ML methods such as LightGBM (ke2017lightgbm) and XGBoost (chen2016xgboost) demonstrated superior performance in capturing complex financial patterns, while DL models including LSTM (hochreiter1997long) and Transformer (vaswani2017attention) architectures have revolutionized temporal modeling. Recent advances in financial forecasting include TimesNet (wu2022timesnet), DLinear (zeng2023are), and Timexer (liu2024timexer), which leverage attention mechanisms for improved prediction accuracy. In algorithmic trading, RL methods such as PPO (schulman2017proximal) and DQN (mnih2015human) have enabled adaptive trading strategies (jiang2017deep; liu2020finrl), with recent work exploring transformer-based approaches (qin2024earnhft). Portfolio management has similarly progressed from classical mean-variance optimization (markowitz1952portfolio) to data-driven approaches using ML (heaton2017deep) and RL (zhang2024reinforcement; kim2025semi). Despite these advances, existing approaches often focus on individual tasks, lacking unified frameworks for comprehensive financial AI development.
@@ 2.2. LLMs and LLM Agents in Financial AI
The integration of LLMs into financial decision-making has followed two main paths. The traditional approach uses pre-training and supervised fine-tuning (SFT), with models like FinBERT (araci2019finbert) and FLANG (shah2022flang) excelling in financial text understanding. Recent progress includes FinGPT (liu2023fingpt), BloombergGPT (wu2023bloomberggpt) with domain-adapted tokenization, and FinQA (chen2021finqa) featuring numerical reasoning. A newer direction combines pre-training with reinforcement learning (RL) to enhance reasoning, as shown by Fin-R1 (liu2025fin) and Fino1 (qian2025fino1). Despite their strengths, LLM-based methods often lack sequential decision-making, are computationally expensive, and struggle with non-stationary markets.
To overcome the limitations of traditional LLMs in sequential decision-making, researchers have introduced agentic mechanisms such as memory, tool use, and self-reflection. Early systems like FinMem (yu2024finmem) and FinRobot (zhou2024finrobot) rely on text-only LLMs enhanced with layered memory, profiling, and chain-of-thought reasoning to improve single-asset trading performance. Building on this, FinAgent (zhang2024multimodal) represents the first multimodal trading agent, which jointly processes news, structured price data, and K-line chart visuals via a tool-augmented dual-reflection architecture, achieving significant improvements in profit. Recent efforts shift toward collaborative multi-agent systems: FinCon (yu2024fincon) adopts a manager–analyst hierarchy with verbal reinforcement, while TradingAgents (xiao2024tradingagents) models a full trading firm with specialized agents coordinating to enhance returns and risk control. Beyond these architectural innovations, recent work has also explored using RL to finetune LLMs Agents (feng2025group; agent-r1), enabling dynamic exploration and reasoning capabilities in real-world environments.
@@ 2.3. Financial AI Platforms
The development of financial AI platforms has been driven by the need to provide standardized environments for research and deployment. Qlib (yang2020qlib) provides a quantitative investment platform that integrates traditional ML pipelines with financial data processing, offering tools for feature engineering, model training, and portfolio optimization.TradeMaster (sun2023trademaster) offers a comprehensive RL-based trading framework with modular components for strategy development, backtesting, and evaluation across multiple financial markets. FinRL-Meta (liu2022finrl) focuses on reinforcement learning environments and benchmarks, providing standardized market simulators and evaluation metrics for RL-based trading strategies. These platforms have significantly contributed to the advancement of financial AI research by providing researchers and practitioners with accessible tools and standardized evaluation frameworks. However, most existing platforms primarily focus on traditional ML, DL, and RL methods, with limited support for LLMs and LLMs Agents, and lack comprehensive multimodal data integration capabilities. This motivates our proposal of FinWorld, a comprehensive framework that covers most critical financial models and tasks.
@@ 3.5. Reinforcement Learning for LLMs
Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) is a policy optimization algorithm designed to efficiently train LLMs in both language and environment-interactive settings. In the context of financial AI, GRPO can be leveraged to endow LLMs with either general financial knowledge and reasoning ability, and specialized trading skills. For example, when applied to financial document analysis or financial question answering, GRPO enables the LLM to align with environment feedback and develop robust understanding of domain-specific content. In these tasks, the LLM is trained to produce high-quality financial responses, with group-level reward normalization providing stable learning signals for fine-tuning.
@@ 4.2. Configuration Layer
The configuration layer of FinWorld is built on mmengine (mmengine2022) and provides a unified and extensible system using dictionaries. It centralizes all experimental settings, including datasets, models, training, and evaluation, in a readable and modular format. Support for configuration inheritance and overrides ensures reproducibility, flexibility, and collaborative development. In the meantime, FinWorld uses a registry mechanism to manage core components such as the dataset and the environment. Each class is registered by its type and instantiated from the configuration, allowing for flexible component management and a decoupled system architecture.
@@ 4.3. Dataset Layer
Downloader Module. This module consists of two main components: i) Market Data Downloader, which standardizes access to heterogeneous financial market data sources, such as FMP (fmp2025) and Alpaca (alpaca2025), supporting multiple time resolutions (e.g., daily, minute-level) and data types (e.g., OHLCV, news); and ii) LLM Reasoning Dataset Downloader, which provides access to financial reasoning datasets including FinQA, professional certification materials (e.g., ACCA, CFA), and other financial knowledge bases. The unified downloader abstraction streamlines data acquisition from various APIs and databases, ensuring consistency and scalability across research tasks.
Processor Module. Building on the standardized data access, this module enables users to configure essential preprocessing steps, such as factor computation (e.g., alpha158 (yang2020qlib)), feature selection, and normalization, LLM reasoning data processing, and other preprocessing steps according to specific research needs. The processor abstraction facilitates reproducible and customizable data transformations for downstream modeling.
@@ 4.4. Model Layer
DL Models. This module organizes neural architectures into components, layers, and model networks. Components include data embedding for financial time series (wu2022timesnet), patch embedding (dosovitskiy2020image), position encodings, attention, transformer blocks, and activation functions. Layers compose these into an embedding layer, an encoder, and a decoder with consistent interfaces. Model networks such as Autoformer, VAE, and GPT-style decoders are instantiated by wiring the layers under a common specification, independent of training and inference.
@@
LLMs Models. This module centralizes access to both proprietary and open‑source LLMs via a unified interface. It supports seamless switching among commercial providers (e.g., GPT‑4.1, Claude‑4‑Sonnet, Gemini‑2.5‑Pro) and efficient local models (e.g., Qwen2.5, Qwen3), with standardized controls for context length, sampling, and cost/latency tracking. Built‑in tools include function calling, tool use, and retrieval‑augmented generation for financial documents. The module integrates naturally with agents and downstream models, enabling document understanding, information extraction, instruction following.
@@ 5.2. Time Series Forecasting
Methods. We evaluate several ML-based and DL-based time series forecasting models, including: i) ML-based methods: LightGBM (yang2020qlib), XGBoost (yang2020qlib); ii) DL-based methods: Autoformer (wu2021autoformer), Crossformer (zhang2023crossformer), ETSformer (woo2022etsformer), DLinear (zeng2023transformers), TimesNet (wu2022timesnet), PatchTST (nie2022time), TimeMixer (wang2024timemixer), and TimeXer (wang2024timexer). These models are selected based on their effectiveness in capturing complex patterns and their ability to handle high-dimensional data.
@@ 5.3. Algorithmic Trading
Methods. We systematically evaluate several Rule-based, ML-based, DL-based, and RL-based methods, including i) Rule-based methods: BUY&HOLD, MACD; ii) ML-based Methods: LightGBM, XGBoost; iii) DL-based Methods: Transformer (vaswani2017attention), LSTM, DLinear (zeng2023transformers); iv) RL-based Methods: PPO (schulman2017proximal), SAC (haarnoja2018soft).
@@ 5.4. Portfolio Management
Methods. We evaluate ML-based, DL-based, and RL-based methods. Since we can apply the Top-$k$ Dropout Strategy  (yang2020qlib) after any ML&DL-based forecasting model to construct a portfolio, our ML&DL-based methods include all ML&DL-based forecasting models. Specifically, we consider: i) Rule-based methods: BUY&HOLD; ii) ML-based Methods: LightGBM, XGBoost; iii) DL-based Methods: Autoformer (wu2021autoformer), Crossformer (zhang2023crossformer), ETSformer (woo2022etsformer), DLinear (zeng2023transformers), TimesNet (wu2022timesnet), PatchTST (nie2022time), TimeMixer (wang2024timemixer), and TimeXer (wang2024timexer); iv) RL-based Methods: PPO (schulman2017proximal), SAC (haarnoja2018soft).
@@ 5.5. LLMs Applications
Methods. For financial reasoning, we evaluate our proposed models FinReasoner against open-source LLMs on the test dataset, including DeepSeek-R1 (guo2025deepseek), Qwen3-8B (yang2025qwen3), Fin-R1-7B (guo2025deepseek), and Qwen2.5-7B-Instruct (qwen2025qwen25technicalreport). For trading abilities, we evaluate our own proposed FinReasoner with DeepSeek-R1 (guo2025deepseek), Qwen3-8B (yang2025qwen3), Fin-R1-7B (guo2025deepseek), Qwen2.5-7B-Instruct (qwen2025qwen25technicalreport), GPT-4.1, Claude-4-Sonnet, all serving as the backbone LLMs within FinAgent framework.
@@ A.1. Market Dataset
The data processing pipeline includes comprehensive preprocessing steps to ensure data quality and usability. For price data, we implement filtering procedures to remove non-trading days and invalid data points, followed by the calculation of technical indicators using the Alpha158 (yang2020qlib) (Table 5) framework, which provides a solid foundation for deep learning modeling. For news text data, we employ a two-stage processing approach: first, we use web scraping scripts to extract raw content from original news links, then we utilize Qwen3-32B to generate concise summaries, enabling efficient text analysis and feature extraction for financial AI applications.
@@
kmid (closeopen)/close(close-open)/close K-Mid price ratio
kmid2 (closeopen)/(highlow)(close-open)/(high-low) Normalized K-Mid price ratio
klen (highlow)/open(high-low)/open K-Length price ratio
kup (highmax(open,close))/open(high-max(open,close))/open K-Up price ratio
kup2 (highmax(open,close))/(highlow)(high-max(open,close))/(high-low) Normalized K-Up price ratio
klow (min(open,close)low)/open(min(open,close)-low)/open K-Low price ratio
klow2 (min(open,close)low)/(highlow)(min(open,close)-low)/(high-low) Normalized K-Low price ratio
@@
roc_ww $close.shift(w)/close$ Rate of change over $w$ periods
ma_ww $close.rolling(w).mean()/close$ Moving average ratio over $w$ periods
std_ww $close.rolling(w).std()/close$ Standard deviation ratio over $w$ periods
beta_ww $(close.shift(w)-close)/(w\times close)$ Beta coefficient over $w$ periods
max_ww $close.rolling(w).max()/close$ Maximum price ratio over $w$ periods
min_ww $close.rolling(w).min()/close$ Minimum price ratio over $w$ periods
qtlu_ww (closeclose.rolling(w).quantile(0.8))/close(close-close.rolling(w).quantile(0.8))/close Upper quantile ratio over $w$ periods
qtld_ww (closeclose.rolling(w).quantile(0.2))/close(close-close.rolling(w).quantile(0.2))/close Lower quantile ratio over $w$ periods
rank_ww close.rolling(w).rank(pct=True).iloc[1]/wclose.rolling(w).rank(pct=True).iloc[-1]/w Price rank percentile over $w$ periods
@@
imax_ww $high.rolling(w).argmax()/w$ Position of maximum high over $w$ periods
imin_ww $low.rolling(w).argmin()/w$ Position of minimum low over $w$ periods
imxd_ww (high.rolling(w).argmax()low.rolling(w).argmin())/w(high.rolling(w).argmax()-low.rolling(w).argmin())/w Distance between max/min positions
@@
rsv_ww (closemin(low,close.shift(w)))/(max(high,close.shift(w))min(low,close.shift(w)))(close-min(low,close.shift(w)))/(max(high,close.shift(w))-min(low,close.shift(w))) RSV (Raw Stochastic Value)
cntp_ww $ret1.gt(0).rolling(w).sum()/w$ Count of positive returns over $w$ periods
cntn_ww $ret1.lt(0).rolling(w).sum()/w$ Count of negative returns over $w$ periods
cntd_ww cntp_wcntn_wcntp\_w-cntn\_w Difference between positive and negative return counts
@@
corr_ww $close.rolling(w).corr(\log(volume+1).rolling(w))$ Correlation between close and log volume
cord_ww $(close/close.shift(1)).rolling(w).corr(\log(volume/volume.shift(1)+1).rolling(w))$ Normalized correlation
@@
sump_ww pos_ret1.rolling(w).sum()/abs_ret1.rolling(w).sum()pos\_ret1.rolling(w).sum()/abs\_ret1.rolling(w).sum() Sum of positive returns ratio
sumn_ww 1sump_w1-sump\_w Sum of negative returns ratio
sumd_ww $2\times sump\_w-1$ Difference between positive and negative return sums
@@
vma_ww $volume.rolling(w).mean()/volume$ Volume moving average ratio
vstd_ww $volume.rolling(w).std()/volume$ Volume standard deviation ratio
wvma_ww |close/close.shift(1)1|.rolling(w).std()/|close/close.shift(1)1|.rolling(w).mean()|close/close.shift(1)-1|.rolling(w).std()/|close/close.shift(1)-1|.rolling(w).mean() Weighted volume moving average
vsump_ww pos_vchg1.rolling(w).sum()/abs_vchg1.rolling(w).sum()pos\_vchg1.rolling(w).sum()/abs\_vchg1.rolling(w).sum() Sum of positive volume changes ratio
vsumn_ww 1vsump_w1-vsump\_w Sum of negative volume changes ratio
vsumd_ww $2\times vsump\_w-1$ Difference between positive and negative volume change sums